OpenStudy (anonymous):

Can someone explain to me how to do this? determine whether y varies directly with x. if so, find the constant of variation k and write the equation. x y 5 20 15 60 45 180 135 540

OpenStudy (cwrw238):

y = kx is direct variation where k is a constant

OpenStudy (cwrw238):

k >= 1 as x increases y will also increase

OpenStudy (cwrw238):

is this the case with your table?

OpenStudy (anonymous):

yeah!

OpenStudy (cwrw238):

right now can you see what k is?

OpenStudy (cwrw238):

5 * k = 20 15 * k = 60 do you see the pattern?

OpenStudy (anonymous):

yeah

OpenStudy (cwrw238):

k = 20/5 = 4 equation is y = 4x

OpenStudy (anonymous):

thanks!!